1.How could a project manager adjust the cost of capital (i.e., appropriate discount rate) to increase the likelihood of having his/her project accepted? Is this ethical or financially sound?

2.The company’s 100,000 shares of preferred stock pay a $3 annual dividend, and sell for $30 per share. Identify the cost of preferred shares?

3.The company’s 500,000 shares of common stock sell for $25 per share and have a beta of 1.5. The risk free rate is 4%, and the market risk premium is 8%.

Identify the cost of equity.

4. Assuming a 40% tax rate, what is the company’s WACC?
Answer #1

1. A manager could assume that the project is less risky than the typical firm project and therefore apply a lower discount rate, which would increase the NPV. This illustrates the importance of sensitivity analysis in evaluating proposed projects .

This is not ethically sound, a manger should project the true and real picture of a project, to prevent accepting risky projects which could lead to wealth destruction.

2. The cost of preferred shares is :

Kp = Dp/ price of preferred shares

= $3/$30

= 10%

Therefore, the cost of preferred capital is 10%.

3. the cost of equity is :

Re = Rf + beta (Rm - Rf)

= 4 + 1.5* 8

=16%

4. the WACC is : weight of equity * cost of equity + weight of preference shares*cost of preference shares

value of equity =$125,00,000

value of preference shares =$3,00,0000

total value = $15,500,000

= 125,00,000/15,500,000* 0.16 + 3,000,000/15,500,000* 0.1

=0.806 *0.16 + 0.1935*0.1

=0.12896 + 0.01935

=14.83%